# Distributional Limits for Eigenvalues of Graphon Kernel Matrices

### v4

@@ 1 Introduction
Our proof combines three probabilistic ingredients. First, a second-order Rayleigh-Schrödinger expansion for simple eigenvalues Kato and Katåo (1966) reduces the problem to the analysis of a centered Rayleigh quotient. Second, concentration of kernel matrices under boundedness assumptions yields operator-norm control of the empirical perturbation. Third, a Hoeffding decomposition for the resulting symmetric statistic Serfling (2009) isolates the linear and degenerate components, revealing the mechanism behind the Gaussian versus weighted chi-square regimes. This synthesis demonstrates that the fluctuation dichotomy persists well beyond the smooth settings considered in Chatterjee and Huang (2025), and that the limiting distributions are determined by spectral structure and degeneracy rather than regularity of the kernel.
@@ 6 Proof of the main theorem
Moreover on this high-probability event, $\eta_{n}-\alpha_{n}\asymp n$, $\beta_{n}-\eta_{n}\asymp n$. The Kato-Temple inequality for self-adjoint matrices (see, e.g., Kato Kato and Katåo (1966)) then yields
@@ Acknowledgments
Conditioning on $U_{i}$, the summands are independent in $j$ and
@@
Moreover, since $|W|\leq\|W\|_{\infty}$ and $\mathbb{E}[\varphi_{r}(U)^{2}]=1$,
$\operatorname{Var}\!\left(W(U_{i},U_{j})\varphi_{r}(U_{j})\mid U_{i}\right)\leq\|W\|_{\infty}^{2}\,\mathbb{E}[\varphi_{r}(U)^{2}]=\|W\|_{\infty}^{2}.$
Hence
$\mathbb{E}[\Delta_{i}^{2}\mid U_{i}]\leq\frac{C}{n}\quad\text{for a constant }C>0,$
and therefore
$\mathbb{E}\left\|\frac{\mathbf{K}_{n}}{n-1}\tilde{\mathbf{u}}_{r}-\lambda_{r}\tilde{\mathbf{u}}_{r}\right\|_{2}^{2}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\Delta_{i}^{2}]\leq\frac{C}{n}.$
@@
Under Assumption 1, let $T_{W}$ denote the integral operator with eigenvalues $(\lambda_{k})_{k\geq 1}$ satisfying the eigengap $\gamma_{r}=\inf_{k\neq r}|\lambda_{r}-\lambda_{k}|>0$. Let $\mathbf{K}_{n}$ be the kernel matrix defined in Lemma 10, and denote its ordered eigenvalues by $\lambda_{1}(\mathbf{K}_{n})\geq\lambda_{2}(\mathbf{K}_{n})\geq\cdots\geq\lambda_{n}(\mathbf{K}_{n})$. Then there exists a constant $c>0$ such that
@@
From Lemma 10,
$\Big\|\frac{1}{n-1}\mathbf{K}_{n}-T_{W}\Big\|_{\mathrm{op}}=O_{p}(n^{-1/2}).$
By Weyl’s inequality for self-adjoint matrices,
$\Big|\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}-\lambda_{k}(T_{W})\Big|\leq\Big\|\frac{1}{n-1}\mathbf{K}_{n}-T_{W}\Big\|_{\mathrm{op}}.$
Hence, with probability 1o(1)1-o(1),
$\min_{k\neq r}\Big|\frac{\lambda_{r}(\mathbf{K}_{n})}{n-1}-\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}\Big|\geq\frac{\gamma_{r}}{2}.$
Multiplying both sides by (n1)(n-1) gives
$\min_{k\neq r}|\lambda_{r}(\mathbf{K}_{n})-\lambda_{k}(\mathbf{K}_{n})|\geq cn$
for some $c=\gamma_{r}/4$, completing the proof. ∎
@@
Suppose there exist real numbers $\alpha<\eta<\beta$ such that
$\mathrm{spec}(\mathbf{A})\cap(\alpha,\beta)=\{\lambda\}$
consists of exactly one eigenvalue $\lambda$ of $\mathbf{A}$. Then
@@
In particular, if $\eta-\alpha$ and $\beta-\eta$ are of order $n$ and $\|\mathbf{r}\|_{2}=o(\sqrt{n})$, then
$|\lambda-\eta|=o_{p}(1).$
@@
This is a finite-dimensional specialization of the Kato-Temple inequality for self-adjoint operators; see  Kato and Katåo (1966, Theorems V.4.10-V.4.11). ∎
@@
so that $\mathbf{Q}$ spans the orthogonal complement of $\mathbf{u}_{r}$ and $\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I}_{n-1},\,\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{n}-\mathbf{u}_{r}\mathbf{u}_{r}^{\top}.$ Every $\mathbf{x}\in\mathbb{R}^{n}$ admits a unique decomposition
@@
$\bigl(\mathbf{Q}^{\top}(\mathbf{M}-\lambda_{r}\mathbf{I})\mathbf{Q}+\mathbf{Q}^{\top}(\mathbf{E}-\alpha\mathbf{I})\mathbf{Q}\bigr)\mathbf{z}=-\xi\,\mathbf{Q}^{\top}\mathbf{E}\mathbf{u}_{r}.$
Set
@@
By the eigengap assumption, the eigenvalues of $\mathbf{L}_{0}$ are $\lambda_{k}(\mathbf{M})-\lambda_{r}$ for $k\neq r$, so $\|\mathbf{L}_{0}^{-1}\|_{\mathrm{op}}\leq 1/\gamma_{r}.$ By Weyl’s inequality, $|\alpha|\leq\|\mathbf{E}\|_{\mathrm{op}}$, hence
$\|\mathbf{\Delta}\|_{\mathrm{op}}\leq\|\mathbf{Q}^{\top}\mathbf{E}\mathbf{Q}\|_{\mathrm{op}}+|\alpha|\leq 2\|\mathbf{E}\|_{\mathrm{op}}<\gamma_{r}.$
Thus $\mathbf{L}_{0}+\mathbf{\Delta}$ is invertible and
@@
$\alpha\xi=\mathbf{u}_{r}^{\top}\mathbf{E}(\mathbf{u}_{r}\xi+\mathbf{Q}\mathbf{z})=\xi\,\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{u}_{r}+\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{Q}\mathbf{z}.$
If $\xi=0$, then $\|\mathbf{z}\|=1$ and the equation above gives $|\alpha|\leq\|\mathbf{E}\|_{\mathrm{op}}$, which is consistent with the claimed bound. Otherwise, divide by $\xi$ to get
$\alpha=\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{u}_{r}+\frac{1}{\xi}\,\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{Q}\mathbf{z}.$
Substitute the expression for $\mathbf{z}$:
@@
Thus $\widetilde{\lambda}_{r}-\lambda_{r}=\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{u}_{r}+R$ with
$R:=-\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{Q}(\mathbf{L}_{0}+\mathbf{\Delta})^{-1}\mathbf{Q}^{\top}\mathbf{E}\mathbf{u}_{r}.$
The norm bound
$|R|\leq\|\mathbf{E}\|_{\mathrm{op}}\cdot\|(\mathbf{L}_{0}+\mathbf{\Delta})^{-1}\|_{\mathrm{op}}\cdot\|\mathbf{E}\|_{\mathrm{op}}\leq\frac{2\|\mathbf{E}\|_{\mathrm{op}}^{2}}{\gamma_{r}}$
follows directly. ∎
@@
$V_{r,n}:=\frac{1}{n}\sum_{i=1}^{n}\bigl(\varphi_{r}(U_{i})^{2}-1\bigr),\qquad s_{n}:=\sum_{i=1}^{n}\varphi_{r}(U_{i})^{2}=n\bigl(1+V_{r,n}\bigr).$
Then
$\frac{1}{s_{n}}=\frac{1}{n}\Bigl(1-V_{r,n}+O_{p}(n^{-1})\Bigr),\qquad\frac{1}{\sqrt{s_{n}}}=\frac{1}{\sqrt{n}}\Bigl(1-\tfrac{1}{2}V_{r,n}+O_{p}(n^{-1})\Bigr).$
@@
By the law of large numbers applied to the i.i.d. variables $\varphi_{r}(U_{i})^{2}$ with $\mathbb{E}[\varphi_{r}(U)^{2}]=1$ and $\text{Var}(\varphi_{r}(U)^{2})=\sigma_{r}^{2}<\infty$, we have $V_{r,n}=O_{p}(n^{-1/2})$.
For the first expansion, note that $s_{n}=n(1+V_{r,n})$, hence
@@
Using the Taylor expansion $(1+x)^{-1}=1-x+x^{2}-x^{3}+\cdots$ for $|x|<1$, we obtain
$(1+V_{r,n})^{-1}=1-V_{r,n}+V_{r,n}^{2}+O(V_{r,n}^{3}).$
Since $V_{r,n}=O_{p}(n^{-1/2})$, it follows that $V_{r,n}^{2}=O_{p}(n^{-1})$ and $V_{r,n}^{3}=O_{p}(n^{-3/2})$. Therefore,
$\frac{1}{s_{n}}=\frac{1}{n}\bigl(1-V_{r,n}+O_{p}(n^{-1})\bigr).$
For the second expansion, we use the Taylor series $(1+x)^{-1/2}=1-\tfrac{1}{2}x+\tfrac{3}{8}x^{2}+O(x^{3})$, which yields
$(1+V_{r,n})^{-1/2}=1-\tfrac{1}{2}V_{r,n}+O(V_{r,n}^{2}).$
Again, since $V_{r,n}^{2}=O_{p}(n^{-1})$, we conclude that
$\frac{1}{\sqrt{s_{n}}}=\frac{1}{\sqrt{n}}\bigl(1-\tfrac{1}{2}V_{r,n}+O_{p}(n^{-1})\bigr).$

### v5

# arXiv:2601.04584v5[math.PR] 28 Feb 2026

@@ 1 Introduction
Our proof combines three probabilistic ingredients. First, a second-order Rayleigh-Schrödinger expansion for simple eigenvalues (Kato and Katåo (1966), Lemma  12) reduces the problem to the analysis of a centered Rayleigh quotient. Second, concentration of kernel matrices under boundedness assumptions yields operator-norm control of the empirical perturbation. Third, a Hoeffding decomposition for the resulting symmetric statistic Serfling (2009) isolates the linear and degenerate components, revealing the mechanism behind the Gaussian versus weighted chi-square regimes. This synthesis demonstrates that the fluctuation dichotomy persists well beyond the smooth settings considered in Chatterjee and Huang (2025), and that the limiting distributions are determined by spectral structure and degeneracy rather than regularity of the kernel.
@@ 6 Proof of the main theorem
Moreover on this high-probability event, $\eta_{n}-\alpha_{n}\asymp n$, $\beta_{n}-\eta_{n}\asymp n$. The Kato-Temple inequality for self-adjoint matrices (see, e.g., Kato Kato and Katåo (1966) or 9) then yields
@@ Acknowledgments
Then $\mathbb{E}[\Delta_{i}\mid U_{i}]=0$. Conditioning on $U_{i}$, the summands are independent in $j$ and
@@
Moreover,
$\operatorname{Var}\!\left(W(U_{i},U_{j})\varphi_{r}(U_{j})\mid U_{i}\right)\leq\mathbb{E}\!\left[W(U_{i},U_{j})^{2}\varphi_{r}(U_{j})^{2}\mid U_{i}\right]\leq\|W\|_{\infty}^{2}\,\mathbb{E}[\varphi_{r}(U)^{2}]=\|W\|_{\infty}^{2}.$
Hence, conditional on $U_{i}$,
$\mathbb{E}[\Delta_{i}^{2}\mid U_{i}]=\operatorname{Var}(\Delta_{i}\mid U_{i})=\frac{1}{(n-1)^{2}}\sum_{j\neq i}\operatorname{Var}\!\left(W(U_{i},U_{j})\varphi_{r}(U_{j})\mid U_{i}\right)\leq\frac{C}{n-1},$
for a constant $C>0$. Therefore,
$\mathbb{E}\left\|\frac{\mathbf{K}_{n}}{n-1}\tilde{\mathbf{u}}_{r}-\lambda_{r}\tilde{\mathbf{u}}_{r}\right\|_{2}^{2}=\sum_{i=1}^{n}\mathbb{E}\left[\left(\frac{1}{\sqrt{n}}\Delta_{i}\right)^{2}\right]=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\Delta_{i}^{2}]\leq\frac{C}{n}.$
@@
Under Assumption 1, let $T_{W}$ denote the integral operator with eigenvalues $(\lambda_{k})_{k\geq 1}$, and assume that $\lambda_{r}$ is simple and isolated:
$\gamma_{r}:=\min_{k\neq r}|\lambda_{r}-\lambda_{k}|>0.$
Let $\mathbf{K}_{n}$ be the kernel matrix and set
$\mathbf{K}_{n}^{\star}:=\mathbb{E}\!\left[\mathbf{K}_{n}\,\middle|\,U_{1},\dots,U_{n}\right].$
Assume further that
$\max_{1\leq k\leq n}\left|\frac{\lambda_{k}(\mathbf{K}_{n}^{\star})}{n-1}-\lambda_{k}\right|=o_{p}(1).$
Then there exists a constant $c>0$ such that
@@
By Lemma 10,
$\Big\|\frac{1}{n-1}\mathbf{K}_{n}-\frac{1}{n-1}\mathbf{K}_{n}^{\star}\Big\|_{\mathrm{op}}=O_{p}(n^{-1/2}).$
By Weyl’s inequality for self-adjoint matrices, for each $k\in[n]$,
$\Big|\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}-\frac{\lambda_{k}(\mathbf{K}_{n}^{\star})}{n-1}\Big|\leq\Big\|\frac{1}{n-1}\mathbf{K}_{n}-\frac{1}{n-1}\mathbf{K}_{n}^{\star}\Big\|_{\mathrm{op}}.$
Combining with the assumption $\max_{1\leq k\leq n}\left|\lambda_{k}(\mathbf{K}_{n}^{\star})/(n-1)-\lambda_{k}\right|=o_{p}(1)$, we obtain that with probability 1o(1)1-o(1),
$\Big|\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}-\lambda_{k}\Big|\leq\frac{\gamma_{r}}{4}\qquad\text{for all }k\in[n].$
Therefore, on this event,
$\min_{k\neq r}\Big|\frac{\lambda_{r}(\mathbf{K}_{n})}{n-1}-\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}\Big|\geq\min_{k\neq r}|\lambda_{r}-\lambda_{k}|-2\cdot\frac{\gamma_{r}}{4}=\frac{\gamma_{r}}{2}.$
Multiplying by $(n-1)\geq n/2$ for large $n$ yields
$\min_{k\neq r}|\lambda_{r}(\mathbf{K}_{n})-\lambda_{k}(\mathbf{K}_{n})|\geq\frac{\gamma_{r}}{4}\,n$
with probability 1o(1)1-o(1), completing the proof. ∎
@@
Suppose there exist real numbers $\alpha<\eta<\beta$ such that the open interval $(\alpha,\beta)$ contains exactly one eigenvalue $\lambda$ of $\mathbf{A}$ (counting multiplicity). Then
@@
In particular, if $\eta-\alpha$ and $\beta-\eta$ are of order $n$ and $\|\mathbf{r}\|_{2}=o(\sqrt{n})$, then $|\lambda-\eta|=o(1)$.
@@
This is a finite-dimensional specialization of the Kato–Temple inequality. Specifically, equation (10) in Kato (1949) states that if the interval $(\alpha,\beta)$ contains no point of the spectrum except at most one (non-degenerate) eigenvalue and
$\varepsilon^{2}<(\eta-\alpha)(\beta-\eta),$
then
$\eta-\frac{\varepsilon^{2}}{\beta-\eta}\leq\lambda\leq\eta+\frac{\varepsilon^{2}}{\eta-\alpha}.$
Identifying $\varepsilon=\|\mathbf{r}\|_{2}$ yields the stated bounds. ∎
@@
so that $\mathbf{Q}$ spans the orthogonal complement of $\mathbf{u}_{r}$ and $\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I}_{n-1},\,\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{n}-\mathbf{u}_{r}\mathbf{u}_{r}^{\top}.$
Every $\mathbf{x}\in\mathbb{R}^{n}$ admits a unique decomposition
@@
$\bigl(\mathbf{L}_{0}+\mathbf{\Delta}\bigr)\mathbf{z}=-\xi\,\mathbf{Q}^{\top}\mathbf{E}\mathbf{u}_{r},$
where
@@
By the eigengap assumption, $\|\mathbf{L}_{0}^{-1}\|_{\mathrm{op}}\leq 1/\gamma_{r}$. Under $\|\mathbf{E}\|_{\mathrm{op}}<\gamma_{r}/2$, $\mathbf{L}_{0}+\mathbf{\Delta}$ is invertible and
@@
$\alpha\xi=\xi\,\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{u}_{r}+\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{Q}\mathbf{z}.$
If $\xi\neq 0$, divide by $\xi$ and substitute $\mathbf{z}$:
@@
We now expand the inverse via a resolvent identity:
$(\mathbf{L}_{0}+\mathbf{\Delta})^{-1}=\mathbf{L}_{0}^{-1}-\mathbf{L}_{0}^{-1}\mathbf{\Delta}\mathbf{L}_{0}^{-1}+\mathbf{R}_{3},$
where $\|\mathbf{R}_{3}\|_{\mathrm{op}}\leq C\|\mathbf{E}\|_{\mathrm{op}}^{2}/\gamma_{r}^{3}.$
Substituting gives
$\alpha=\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{u}_{r}-\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{Q}\mathbf{L}_{0}^{-1}\mathbf{Q}^{\top}\mathbf{E}\mathbf{u}_{r}+R_{3},$
with $|R_{3}|\leq C\|\mathbf{E}\|_{\mathrm{op}}^{3}/\gamma_{r}^{2}.$
Since
$\mathbf{Q}\mathbf{L}_{0}^{-1}\mathbf{Q}^{\top}=\sum_{k\neq r}\frac{1}{\lambda_{k}-\lambda_{r}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top},$
we obtain the explicit second-order Rayleigh–Schrödinger term
$\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{Q}\mathbf{L}_{0}^{-1}\mathbf{Q}^{\top}\mathbf{E}\mathbf{u}_{r}=\sum_{k\neq r}\frac{(\mathbf{u}_{k}^{\top}\mathbf{E}\mathbf{u}_{r})^{2}}{\lambda_{r}-\lambda_{k}}.$
Therefore,
$\widetilde{\lambda}_{r}-\lambda_{r}=\mathbf{u}_{r}^{\top}\mathbf{E}\mathbf{u}_{r}+\sum_{k\neq r}\frac{(\mathbf{u}_{k}^{\top}\mathbf{E}\mathbf{u}_{r})^{2}}{\lambda_{r}-\lambda_{k}}+R_{3},$
with
$|R_{3}|\leq C\frac{\|\mathbf{E}\|_{\mathrm{op}}^{3}}{\gamma_{r}^{2}}.$
This yields the full second-order Rayleigh–Schrödinger expansion. ∎
@@
$V_{r,n}:=\frac{1}{n}\sum_{i=1}^{n}(\varphi_{r}(U_{i})^{2}-1),\qquad s_{n}:=\sum_{i=1}^{n}\varphi_{r}(U_{i})^{2}=n(1+V_{r,n}).$
Assume $\operatorname{Var}(\varphi_{r}(U)^{2})<\infty$. Then $V_{r,n}=O_{p}(n^{-1/2})$ and
$\frac{1}{s_{n}}=\frac{1}{n}\Bigl(1-V_{r,n}+O_{p}(V_{r,n}^{2})\Bigr),\qquad\frac{1}{\sqrt{s_{n}}}=\frac{1}{\sqrt{n}}\Bigl(1-\tfrac{1}{2}V_{r,n}+O_{p}(V_{r,n}^{2})\Bigr).$
In particular, since $V_{r,n}^{2}=O_{p}(n^{-1})$, both remainders are $O_{p}(n^{-1})$.
@@
Assume $\operatorname{Var}(\varphi_{r}(U)^{2})<\infty$ and recall
$V_{r,n}:=\frac{1}{n}\sum_{i=1}^{n}\bigl(\varphi_{r}(U_{i})^{2}-1\bigr),\qquad s_{n}:=\sum_{i=1}^{n}\varphi_{r}(U_{i})^{2}=n(1+V_{r,n}).$
Since $\mathbb{E}[\varphi_{r}(U)^{2}]=1$ and $\operatorname{Var}(\varphi_{r}(U)^{2})<\infty$, we have
$\mathbb{E}[V_{r,n}]=0,\qquad\operatorname{Var}(V_{r,n})=\frac{1}{n}\operatorname{Var}(\varphi_{r}(U)^{2}),$
hence by Chebyshev’s inequality $V_{r,n}=O_{p}(n^{-1/2})$.
For the first expansion, write
@@
Using the Taylor expansion of $f(x)=(1+x)^{-1}$ at 0,
$(1+x)^{-1}=1-x+x^{2}+R_{1}(x),$
where the remainder satisfies $|R_{1}(x)|\leq C|x|^{3}$ for all $|x|\leq 1/2$ and a universal constant $C>0$. On the event $\{|V_{r,n}|\leq 1/2\}$,
$(1+V_{r,n})^{-1}=1-V_{r,n}+V_{r,n}^{2}+R_{1}(V_{r,n}),\qquad|R_{1}(V_{r,n})|\leq C|V_{r,n}|^{3}.$
Since $V_{r,n}=O_{p}(n^{-1/2})$, we have $V_{r,n}^{2}=O_{p}(n^{-1})$ and $V_{r,n}^{3}=O_{p}(n^{-3/2})$, so
$(1+V_{r,n})^{-1}=1-V_{r,n}+O_{p}(V_{r,n}^{2}).$
Multiplying by $1/n$ yields
$\frac{1}{s_{n}}=\frac{1}{n}\Bigl(1-V_{r,n}+O_{p}(V_{r,n}^{2})\Bigr).$
For the second expansion, similarly
$\frac{1}{\sqrt{s_{n}}}=\frac{1}{\sqrt{n}}(1+V_{r,n})^{-1/2}.$
Using the Taylor expansion of $g(x)=(1+x)^{-1/2}$ at 0,
$(1+x)^{-1/2}=1-\tfrac{1}{2}x+R_{2}(x),$
where $|R_{2}(x)|\leq C^{\prime}x^{2}$ for all $|x|\leq 1/2$ and a universal constant $C^{\prime}>0$. On $\{|V_{r,n}|\leq 1/2\}$ we obtain
$(1+V_{r,n})^{-1/2}=1-\tfrac{1}{2}V_{r,n}+R_{2}(V_{r,n}),\qquad|R_{2}(V_{r,n})|\leq C^{\prime}V_{r,n}^{2},$
hence
$(1+V_{r,n})^{-1/2}=1-\tfrac{1}{2}V_{r,n}+O_{p}(V_{r,n}^{2}).$
Multiplying by $1/\sqrt{n}$ gives
$\frac{1}{\sqrt{s_{n}}}=\frac{1}{\sqrt{n}}\Bigl(1-\tfrac{1}{2}V_{r,n}+O_{p}(V_{r,n}^{2})\Bigr).$
Since $V_{r,n}^{2}=O_{p}(n^{-1})$, both $O_{p}(V_{r,n}^{2})$ remainders are also $O_{p}(n^{-1})$. ∎